the height of a tv transmission tower at any place on the surface of the earth is  245m. the maximum distance up to which  transmission of tower will reach is

Grade:11

2 Answers

Vikas TU
7 years ago
From the transmissin concept of Physics,
get to know the relation b/w the Radii of the earth and the maximum distance,
that is:
d = root(2gh) => root(2*10*245) => 4900 m.
Kuldeep kumar
6 years ago
We have to find DtGivenWe have ht 245m ,ht=245mR is radius of earth (6400km) , R=6400kmDt=√2*R*htDt=√2*6400*0.245kmDt=√12800*0.245kmDt=√3136kmDt=56km